Find an angle which is 25 degree less than 4 times it complement
Answer by algebrahouse.com(1659)

Complementary angles add up to 90°

x = an angle
4x - 25 = its complementary angle

x + 4x - 25 = 90 {added the two angles and set equal to 90}
5x - 25 = 90 {combined like terms}
5x = 115 {added 25 to each side}
x = 23 {divided each side by 5}
4x - 25 = 67 {substituted 23, in for x, into 4x - 25}

23° and 67° are the two complementary angles

For more help from me, visit: www.algebrahouse.com